OpenStudy (ariellesircy):

Find the solution set of each equation. If the equation has no solution write 0/ --- If the equation is an identity state so. w^2-7w=0

OpenStudy (owlcoffee):

In order to solve: \(w^2-7w=0\) we will make use of a proposition called "common factor" which can be called a reverse step from the distributive axiom. Now, common factor is pretty much dividing and multiplying the whole thing by the variable in question, o by associative: \[w^2-7w=0 \iff (w^2-7w)=0\] Notice that there is an axiom called the "Neutral of multiplication" it states that there exists a number which multiplied by any quantity, gives as a result, the quantity it multiplies more mathematically written: \(\exists "1" \in \mathbb{R} / 1.a=a\) so, applying it: \[(w^2-7w)=0 \iff 1.(w^2-7w)=0\] \[\frac{ w }{ w } =1\] So, combining these two statements, we get: \[\frac{ w }{ w }(w^2-7w)=0\] Now it's just a matter of operations: \[w(\frac{ w^2 }{ w }-\frac{ 7w }{ w })=0\] \[\iff w(w-7)=0\] Now, from this last step I'll let you continue, try using hankel's property which is as follows: \[a.b=0 \iff a=0; b=0\]

OpenStudy (ariellesircy):

Tbh, I don't even know what Hankel's Property is. This question just has me completely turned upside down

OpenStudy (owlcoffee):

So, beginning from \[w(w-7)=0\] Hankel's property is also called the "zero product rule" by some teachers, which I do not really support. Anywho, if we apply hankel's peroperty, we will have two cases in order for the equality \(w(w-7)=0\) to be satisfied, this being either \(w=0\) is equal to zero or \(w-7=0\) So, we have now the following separate equalities: \[w=0\] \[w-7=0 \] Which I believe you can solve.

OpenStudy (ariellesircy):

w=7?

OpenStudy (owlcoffee):

Good, that is a solution indeed.

OpenStudy (ariellesircy):

When I put that in my answer system it says it's wrong

OpenStudy (owlcoffee):

Well, it must say it's rong because there is not only one solution, one of them is as you stated w=7. Which is the second solution?